OpenStudy (anonymous):

5x + y = 6 5x + 3y = -4 The y-coordinate of the solution to the system shown is _____.

OpenStudy (anonymous):

You can use either elimination or substitution

OpenStudy (anonymous):

?

OpenStudy (anonymous):

How do you want to solve this equation. If you have no idea, we will use elimination

OpenStudy (anonymous):

Ya lets go with elimination

OpenStudy (anonymous):

So first, you multiply the first equation by -1

OpenStudy (anonymous):

-5x -6

OpenStudy (anonymous):

that gives you -5x-y=-6 5x+3y=-4

OpenStudy (anonymous):

add them together

OpenStudy (anonymous):

when you add them together, you get 2y=-10

OpenStudy (anonymous):

wait wouldnt it be 2y=-11 since 5+6=11?

OpenStudy (anonymous):

No, -5x + 5x = 0. So 0 goes in front. -y+3y=2y so 2y goes next. Then -6-4=-10 so -10 goes in the back. That leaves 2y=-10. Then solve for y

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

how do i solve for y?

OpenStudy (anonymous):

divide by 2

OpenStudy (anonymous):

oh -5?

OpenStudy (anonymous):

yes. Now plug that into the first equation 5x+y=6

OpenStudy (anonymous):

so 5x+-5=6

OpenStudy (anonymous):

Yes, now solve for 5

OpenStudy (anonymous):

i meant solve for x

OpenStudy (anonymous):

divide by 2 again?

OpenStudy (anonymous):

No, first, add 5 to both sides to get 5x=11.

OpenStudy (anonymous):

Then, you divide by 5 to get rid of the 5 in front of the x

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

What did you get for x?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

No, the answer is 11/5

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

The reason why is because 5x=11 all divided by 5 leaves x=11/5

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

So the answers are x=11/5 and y=-5

OpenStudy (anonymous):

Ok thank you so much